OpenStudy (anonymous):

What are the first five terms of the sequence given by the formula an = 1n

geerky42 (geerky42):

\(a_n = 1^n\)?

geerky42 (geerky42):

Exactly what is formula here?

OpenStudy (anonymous):

idk thats the question

geerky42 (geerky42):

What is formula exactly? \(a_n = 1n\) \(a_n = 1^n\) \(a_n = 10n\) OR \(a_n = 10^n\) ?

geerky42 (geerky42):

What do formula read as?

OpenStudy (anonymous):

the second one

geerky42 (geerky42):

OK, well then it should be easy. First five terms are \(a_1,a_2,a_3,a_4,a_5\). So to find \(a_1\), just plug in n=1; \(a_1 = 1^1 = ?\)

geerky42 (geerky42):

Do the same for rest of terms.

OpenStudy (anonymous):

thank you